# MOLA: Enhancing Industrial Process Monitoring Using Multi-Block Orthogonal Long Short-Term Memory Autoencoder

### v1

@@ Abstract
In this work, we introduce MOLA: a Multi-block Orthogonal Long short-term memory Autoencoder paradigm, to conduct accurate, reliable fault detection of industrial processes. To achieve this, MOLA effectively extracts dynamic orthogonal features by introducing an orthogonality-based loss function to constrain the latent space output. This helps eliminate the redundancy in the features identified, thereby improving the overall monitoring performance. On top of this, a multi-block monitoring structure is proposed, which categorizes the process variables into multiple blocks by leveraging expert process knowledge about their associations with the overall process. Each block is associated with its specific Orthogonal Long short-term memory Autoencoder model, whose extracted dynamic orthogonal features are monitored by distance-based Hotelling’s $T^{2}$ statistics and quantile-based cumulative sum (CUSUM) designed for multivariate data streams that are nonparametric, heterogeneous in nature. Compared to having a single model accounting for all process variables, such a multi-block structure improves the overall process monitoring performance significantly, especially for large-scale industrial processes. Finally, we propose an adaptive weight-based Bayesian fusion (W-BF) framework to aggregate all block-wise monitoring statistics into a global statistic that we monitor for faults, with the goal of improving fault detection speed by assigning weights to blocks based on the sequential order where alarms are raised. We demonstrate the efficiency and effectiveness of our MOLA framework by applying it to the Tennessee Eastman Process and comparing the performance with various benchmark methods.
@@ 1 Introduction
Effective, reliable process monitoring is essential to ensuring process safety, improving product quality, and reducing operating cost of industrial systems as they continue to expand in scale and complexity [1]. Nowadays, modern chemical plants are equipped with numerous sensors that are connected to Distributed Control Systems (DCSs), generating vast amount of process data that can be leveraged for data-driven process monitoring in real time [2]. Traditional methods for process monitoring encompass Principal Component Analysis (PCA), Partial Least Squares (PLS), and Independent Component Analysis (ICA), among many others [3]. The idea behind these fault detection methods largely falls in extracting underlying features that characterize process states (e.g., faulty vs. non-faulty) from historical process data and monitoring changes in these extracted features [4]. For instance, PCA leverages linear orthogonal transformations to extract key process features, projecting them in a principal component subspace and a residual subspace, followed by developing a monitoring statistic for each subspace for fault detection [5]. Nevertheless, the relationships among process variables being monitored in modern industrial systems are often highly nonlinear in nature, and conventional linear methods such as PCA face challenges in effectively capturing these nonlinear relationships. To overcome the intrinsic limitations of conventional linear methods, kernel-based methods, such as kernel PCA, have been proposed [6, 7]. While these kernel-based methods can extract nonlinear relationships to some extent, identifying and computing the kernel functions can be time consuming, limiting their capabilities in real world applications that demand fast real-time fault detection [4]. Furthermore, compared to standard PCA, kernel-based methods exhibits greater sensitivity to noise and outliers [8], which deteriorate monitoring accuracy.
Leveraging the recent breakthroughs in deep learning, Artificial Neural Networks (ANNs), through the use of multiple fully-connected layers and nonlinear activation functions, have achieved remarkable successes in extracting complex nonlinear features among process variables [9] for process monitoring. Among prevailing ANN-based methods, Multi-layer Perceptron (MLP), Convolutional Neural Network (CNN), and Recurrent Neural Network (RNN) are some of the most effective and widely-used deep learning architectures for process monitoring [10, 11, 12]. Since traditional ANN-based process monitoring methods are essentially supervised classification methods[13], their monitoring performance relies on the availability of a large amount of labeled data, especially faulty data which are typically limited and hard to acquire in practice[14].
@@
Figure 1: Illustration of of autoencoder structure and feature extraction.
@@
In traditional AE-based process monitoring methods, the reconstruction error is used as the primary objective during training and is monitored for fault detection during deployment [16]. However, this approach does not explicitly make use of the lower-dimensional embeddings which inherently represent the underlying process dynamics. Furthermore, solely minimizing the reconstruction error may introduce redundancy among extracted features (see Figure 1). This is illustrated in Figure 1, where the points shown represent the projections of the original input data onto a two-dimensional latent space after passing through the encoder layer. Here, the red point $F$ denotes the projection of a faulty data sample, whereas the rest represent non-faulty data samples. When the extracted features contain redundancy, the two dimensions of the latent space are not orthogonal to each other. In this case, the projections of $F$ onto $X$- and $Y$-axes will fall in the range of non-faulty data sample projections, thereby making fault detection more challenging. On the other hand, when the extracted features contain no redundancy, the latent variables are orthogonal to one another. In the same example, the projection of $F$ onto $X^{\prime}$-axis now falls outside of the range of non-faulty data sample projections.
@@
Based on this observation, Orthogonal Autoencoder (OAE), which introduces a term characterizing orthogonality of latent space outputs in the loss function, was proposed to extract features that are inherently independent or nonredundant [17]. Cacciarelli et al. then applied the OAE to process monitoring applications by monitoring the Hotelling’s $T^{2}$ statistics of orthogonal latent features [18]. While OAE effectively improves the fault detection performance compared to traditional AE, it still has several limitations. First, actual industrial systems typically involve dynamic behaviors [19], which cannot be captured by existing OAE-based framework. Second, the Hotelling’s $T^{2}$ statistic may not be suited for capturing changes and shifts in the distribution of latent features [20]. It also does not explicitly account for the cumulative effects of process anomalies, which can play a crucial role in detecting certain types of faults in a timely manner [21].
@@
Furthermore, one of the unique characteristics of modern chemical process systems is that they typically comprise multiple heavily integrated (via mass, energy, and information flows) yet relatively autonomous subsystems, each with specific process functions such as raw material processing, reaction, and product separations. These subsystems are further disaggregated into smaller unit operations, such as reactors, heat exchangers, distillation columns. Such structural complexity and hierarchy can pose significant challenges to conventional process monitoring paradigms that rely on a single model to monitor the entire process [22]. As the number of process variables being monitored increases, the number of hyperparameters in the deep learning model increases exponentially, significantly amplifying the training complexity. To address this challenge, a multi-block process monitoring methodology has been proposed for large-scale industrial process systems [23]. The idea is to categorize process variables into various blocks based on the variables’ associations with the overall process and their relevance with other process variables, followed by building a process monitoring model for each block. The entire process will be monitored by integrating these block-wise process monitoring models via a data fusion mechanism [24]. Conventional data fusion techniques, such as Bayesian fusion, are static in nature and treat the monitoring statistics from different blocks equally [25]. However, in reality, process anomalies often stems from one block and propagate/spread to others as time progresses, making conventional data fusion techniques inadequate and less effective.
@@
To address the aforementioned challenges, here we propose a novel process monitoring framework based on Multi-block Orthogonal Long short-term memory Autoencoder (MOLA). As a significant variant of traditional autoencoders, Long Short-Term Memory Autoencoder (LSTM AE) implements the LSTM architecture in both the encoder and decoder layers. This allows the extraction of dynamic process features from the time series data. On top of this, we propose the Orthogonal LSTM Autoencoder (OLAE), which incorporates an orthogonality-based loss function to constrain the latent space output. We also adopt the multi-block monitoring methodology to assign all process variables into several blocks based on process knowledge. And a local OLAE model is developed for each block. To effectively detect anomalies of the orthogonal latent features in each block, in addition to Hotelling’s $T^{2}$ approach, we also incorporate a quantile-based multivariate cumulative sum (CUSUM) process monitoring method [26], a state of the art approach to monitor high-dimensional data streams that are nonparametric (i.e., data streams do not necessarily follow any specific distribution) and heterogeneous (i.e., data streams do not necessarily follow the same distribution) [27]. The use of quantile-based multivariate CUSUM successfully overcomes the barrier of Hotelling’s $T^{2}$ method that overlooks the changes and shifts in the distribution of latent features. Finally, we propose an adaptive weight-based Bayesian fusion (W-BF) framework to effectively aggregate the monitoring results from individual blocks. Our proposed W-BF framework automatically assigns higher weights to blocks based on how early anomalies occur in each block, thereby improving overall fault detection speed and accuracy.
@@
The rest of this paper is organized as follows. Section 2 provides a brief review of LSTM and the quantile-based multivariate CUSUM method, followed by a detailed introduction to the proposed MOLA and the adaptive W-BF frameworks. Section 3 discusses the detailed steps involved in offline training and online monitoring of the proposed process monitoring framework. Next, in Section 4, we showcase the performance of our proposed methodology in the benchmark problem of Tennessee Eastman Process (TEP), demonstrating its outstanding fault detection speed and accuracy compared to benchmark methods. To conclude, we summarize all results and learnings and discuss potential improvements for future research in Section 5.
@@ 2.1 LSTM
The recurrent neural network (RNN) is a class of neural network architectures specifically designed for time series modeling and prediction. This makes RNN-based methods particularly attractive in capturing the dynamic features of industrial process data [28]. However, traditional RNNs are prone to the issues of gradient explosion and gradient vanishing when working with long-term time series [29]. To overcome these issues, LSTM, an advanced RNN that uses “gates” to capture both long-term and short-term memory, are introduced, featuring a unique structural unit at their core [30]. As illustrated in Figure 2, an LSTM unit contains three crucial control gates: the forget gate $f(t)$, the input gate $i(t)$, and the output gate $o(t)$. These gates work collectively to ensure that essential information is consistently retained while the less important information is discarded. Therefore, LSTM achieves superior performance in capturing complex temporal dynamics embedded in chemical process systems. For more detailed information and mathematical descriptions about LSTM, readers are encouraged to refer to [31].
@@ 2.2 OLAE
The LSTM AE is a specialized type of AE that seamlessly integrates LSTM with AE. This hybrid architecture enables efficient encoding and decoding of temporal sequences while capturing long-term dynamic features and dependencies within individual data streams at the same time. Similar to AE, the LSTM AE typically uses the reconstruction error as the loss function for model training. The Mean Squared Error (MSE) is one of the most widely adopted reconstruction error:
@@
Figure 3: Our proposed OLAE architecture consists of an LSTM encoder, an LSTM decoder, and a fully-connected (FC) layer that leverages orthogonality. The FC layer is denoted as $C=\sigma(Wh+b)$, where $h$ and $C$ represent the output of encoder and fully-connected (FC) layer, respectively. Here, $b$ and $\sigma$ are the bias term and nonlinear activation function of the FC layer, respectively.
Therefore, in the OLAE architecture as shown in Figure 3, we define an orthogonality-based loss function in Equation (2) to constrain the latent space output so as to to generate non-redundant orthogonal latent features:
@@
where $W=[w_{1},w_{2},\dots,w_{m}]^{T}$ is the weight matrix of the FC layer. The loss function $\operatorname{Loss}_{\perp}$ consists of two components. The first component ensures that the different weights are orthogonal to one another, whereas the second term ensures that the extracted latent features remain as mutually independent as possible even after undergoing nonlinear transformation ($\sigma$).
@@ 2.3 Monitoring statistics
Overall, we monitor the two-sided statistic $W_{x}(t)=\max(W_{x}^{-}(t),W_{x}^{+}(t))$ for both upward and downward mean shifts. An alarm is raised (i.e., an anomaly is detected) when $W_{x}$ exceeds a threshold $h$ that is related to the pre-specified false alarm rate (e.g., 0.0027 for the typical $3\sigma$-limit). The corresponding stopping time $T$ is thus:
$T=\inf\left\{t>0:\sum_{(x)=1}^{r}W_{(x)}(t)\geq h\right\},$ (8)
where $\sum_{(x)=1}^{r}$ denotes the summation of the largest $r$ local statistics and $W_{(x)}(t)$ is the $x$-th largest local statistic [34]. More information about the theory and application of quantile-based multivariate CUSUM, including detailed derivations of the mathematical formulations agove, can be found in Ye and Liu [26].
@@ 2.4 Adaptive weight-based Bayesian fusion strategy
Here, we describe our multi-block monitoring framework for large-scale, complex industrial systems. Each block is monitoed by two metrics (resp. $T^{2}$ and $W$) using two approaches (resp. Hotelling’s $T^{2}$ and quantile-based CUSUM). We adopt a Bayesian data fusion framework to aggregate the two monitoring results into a single monitoring metric in each block, as well as to aggregate all block-level metrics into a single plant-wide fault index (PFI). Bayesian fusion has shown remarkable robustness and capabilities in integrating information from diverse sources. It leverages prior knowledge and real-time measurements to compute posterior probabilities using the Bayes’ theorem. In this work, we extend the classic Bayesian fusion methodology and propose an adaptive weight-based Bayesian fusion (W-BF) method. The idea is to dynamically adjust fusion weights based on the relative ranking of the current monitoring statistics from each block. Under this framework, the probability of sample $X_{i}^{n}(t)$ of block $n$ and monitoring metric $i$ in normal and fault conditions are given by:
@@
$B^{n}(t)=\sum_{i=1}^{2}\frac{w_{i}^{n}(t)\mathbb{P}_{i}^{n}(X_{i}^{n}(t)|F)\mathbb{P}^{n}_{i}(F|X_{i}^{n}(t))}{\sum_{i=1}^{2}w_{i}^{n}(t)\mathbb{P}_{i}^{n}(X_{i}^{n}(t)|F)},$ (11)
@@
$\operatorname{PFI}(t)=\sum_{n=1}^{N}\frac{w_{n}(t)\mathbb{P}^{n}(B^{n}(t)|F)\mathbb{P}^{n}(F|B^{n}(t))}{{\textstyle\sum_{n=1}^{N}}w_{n}(t)\mathbb{P}^{n}(B^{n}(t)|F)};$ (14)

### v2

@@ Abstract
In this work, we introduce MOLA: a Multi-block Orthogonal Long short-term memory Autoencoder paradigm, to conduct accurate, reliable fault detection of industrial processes. To achieve this, MOLA effectively extracts dynamic orthogonal features by introducing an orthogonality-based loss function to constrain the latent space output. This helps eliminate the redundancy in the features identified, thereby improving the overall monitoring performance. On top of this, a multi-block monitoring structure is proposed, which categorizes the process variables into multiple blocks by leveraging expert process knowledge about their associations with the overall process. Each block is associated with its specific Orthogonal Long short-term memory Autoencoder model, whose extracted dynamic orthogonal features are monitored by distance-based Hotelling’s $T^{2}$ statistics and quantile-based cumulative sum (CUSUM) designed for multivariate data streams that are nonparametric and heterogeneous. Compared to having a single model accounting for all process variables, such a multi-block structure significantly improves overall process monitoring performance, especially for large-scale industrial processes. Finally, we propose an adaptive weight-based Bayesian fusion (W-BF) framework to aggregate all block-wise monitoring statistics into a global statistic that we monitor for faults. Fault detection speed and accuracy are improved by assigning and adjusting weights to blocks based on the sequential order in which alarms are raised. We demonstrate the efficiency and effectiveness of our MOLA framework by applying it to the Tennessee Eastman Process and comparing the performance with various benchmark methods.
@@ 1 Introduction
Effective, reliable process monitoring is essential to ensuring process safety, improving product quality, and reducing operating costs of industrial systems as they continue to expand in scale and complexity [1]. Nowadays, modern chemical plants are equipped with numerous sensors connected to Distributed Control Systems (DCSs), continuously generating massive process data that can be leveraged for data-driven process monitoring in real time [2]. Traditional methods for process monitoring encompass Principal Component Analysis (PCA), Partial Least Squares (PLS), and Independent Component Analysis (ICA), among many others [3]. The idea behind these fault detection methods largely falls in extracting underlying features that characterize process states (e.g., faulty vs. non-faulty) from historical process data and monitoring changes in these extracted features [4]. For instance, PCA leverages linear orthogonal transformations to extract key process features, projecting them in a principal component subspace and a residual subspace, followed by developing a monitoring statistic for each subspace for fault detection [5]. Nevertheless, the relationships among process variables being monitored in modern industrial systems are often highly nonlinear, and conventional linear methods such as PCA face challenges in effectively capturing these nonlinear relationships. To overcome the intrinsic limitations of conventional linear methods, kernel-based methods, such as kernel PCA, have been proposed [6, 7]. While these kernel-based methods can extract nonlinear relationships, identifying and computing the kernel functions can be time-consuming, limiting their capabilities in real-world applications that demand fast real-time fault detection [4]. Furthermore, compared to standard PCA, kernel-based methods exhibit greater sensitivity to noise and outliers [8], potentially deteriorating monitoring accuracy.
Leveraging the recent breakthroughs in deep learning, Artificial Neural Networks (ANNs), which consist of multiple fully connected layers and nonlinear activation functions, have achieved remarkable successes in extracting complex nonlinear features among process variables [9] for process monitoring. Among prevailing ANN-based methods, Multi-layer Perceptron (MLP), Convolutional Neural Network (CNN), and Recurrent Neural Network (RNN) are some of the most effective and widely-used deep learning architectures for process monitoring [10, 11, 12]. Since traditional ANN-based process monitoring methods are essentially supervised classification methods [13], their monitoring performance relies on the availability of a large amount of labeled data, especially faulty data which are typically limited and hard to acquire in practice [14].
@@
Figure 1: Illustration of autoencoder structure and feature extraction.
@@
In traditional AE-based process monitoring methods, the reconstruction error is used as the primary objective during training and is monitored for fault detection during deployment [16]. However, this approach does not explicitly make use of the lower-dimensional embeddings which inherently represent the underlying process dynamics. Furthermore, solely minimizing the reconstruction error may introduce redundancy among extracted features (see Figure 1). This is illustrated in Figure 1, where the points shown represent the projections of the original input data onto a two-dimensional latent space after passing through the encoder layer. Here, the red point $F$ denotes the projection of a faulty data sample, whereas the rest represent non-faulty data samples. When the extracted features contain redundancy, the two dimensions of the latent space are not orthogonal to each other. In this case, the projections of $F$ onto $X$- and $Y$-axes will fall in the range of non-faulty data sample projections, thereby making fault detection more challenging. On the other hand, when the extracted features contain no redundancy, the latent variables are orthogonal to one another. In the same example, the projection of $F$ onto the $X^{\prime}$-axis now falls outside of the range of non-faulty data sample projections.
@@
Based on this observation, Orthogonal Autoencoder (OAE), which introduces a term characterizing orthogonality of latent space outputs in the loss function, was proposed to extract features that are inherently independent or nonredundant [17]. Cacciarelli et al. then applied the OAE to process monitoring applications by monitoring Hotelling’s $T^{2}$ statistics of orthogonal latent features [18]. While OAE effectively improves the fault detection performance compared to traditional AE, it still has several limitations. First, actual industrial systems typically involve dynamic behaviors [19], which cannot be captured by existing OAE-based framework. Second, Hotelling’s $T^{2}$ statistic may not be suited for capturing changes and shifts in the distribution of latent features [20]. It also does not explicitly account for the cumulative effects of process anomalies, which can play a crucial role in detecting certain types of faults promptly [21].
@@
Furthermore, one of the unique characteristics of modern chemical process systems is that they typically comprise multiple heavily integrated (via mass, energy, and information flows) yet relatively autonomous subsystems, each with specific process functions such as raw material processing, reaction, and product separations. These subsystems are further disaggregated into smaller unit operations, such as reactors, heat exchangers, and distillation columns. Such structural complexity and hierarchy can pose significant challenges to conventional process monitoring paradigms that rely on a single model to monitor the entire process [22]. As the number of process variables being monitored increases, the number of hyperparameters in the deep learning model increases exponentially, significantly amplifying the training complexity. To address this challenge, a multi-block process monitoring methodology has been proposed for large-scale industrial process systems [23]. The idea is to categorize process variables into various blocks based on the variables’ associations with the overall process and their relevance with other process variables, followed by building a process monitoring model for each block. The entire process will be monitored by integrating these block-wise process monitoring models via a data fusion mechanism [24]. Conventional data fusion techniques, such as Bayesian fusion, are static in nature and treat the monitoring statistics from different blocks equally [25]. However, in reality, process anomalies often stem from one block and propagate/spread to others as time progresses, making conventional data fusion techniques inadequate and less effective.
@@
To address the aforementioned challenges, we propose a novel process monitoring framework based on Multi-block Orthogonal Long short-term memory Autoencoder (MOLA). As a significant variant of traditional autoencoders, Long Short-Term Memory Autoencoder (LSTM AE) implements the LSTM architecture in both the encoder and decoder layers. This allows the extraction of dynamic process features from the time series data. On top of this, we propose the Orthogonal LSTM Autoencoder (OLAE), which incorporates an orthogonality-based loss function to constrain the latent space output. We also adopt the multi-block monitoring methodology to assign all process variables into several blocks based on process knowledge. A local OLAE model is developed for each block. To effectively detect anomalies of the orthogonal latent features in each block, in addition to Hotelling’s $T^{2}$ approach, we also incorporate a quantile-based multivariate cumulative sum (CUSUM) process monitoring method [26], a state-of-the-art approach to monitor high-dimensional data streams that are nonparametric (i.e., data streams do not necessarily follow any specific distribution) and heterogeneous (i.e., data streams do not necessarily follow the same distribution) [27]. The use of quantile-based multivariate CUSUM successfully overcomes the barrier of Hotelling’s $T^{2}$ method that overlooks the changes and shifts in the distribution of latent features. Finally, we propose an adaptive weight-based Bayesian fusion (W-BF) framework to effectively aggregate the monitoring results from individual blocks. Our proposed W-BF framework automatically assigns higher weights to blocks based on how early anomalies occur in each block, thereby improving overall fault detection speed and accuracy. Among all these technical advancements, the key contributions of this work are:
  1. 1.
    We introduce a novel autoencoder architecture OLAE to extract non-redundant and mutually independent dynamic features. Compared to existing autoencoder designs, OLAE demonstrates superior fault detection performance.
  2. 2.
    We adopt a state-of-the-art quantile-based multivariate CUSUM framework to enable fast, accurate, and robust detection of process anomalies based on the mean shift in the distribution of extracted features.
  3. 3.
    We propose a novel W-BF approach to dynamically adjust the weights assigned to monitoring results from different blocks, which significantly enhances fault detection speed and accuracy.
Dataset from the benchmark Tennessee Eastman Process (TEP) problem are used to evaluate our proposed MOLA framework. Some of the key capabilities of MOLA with respect to other related process monitoring methods are summarized in Table 1.
Table 1: A high-level comparison of key capabilities of different process monitoring methods.
Capability PCA KPCA DPCA AE LSTM AE Block PCA Block LSTM AE MOLA Cross-correlation Dynamic $\times$ $\times$ $\times$ $\times$ Nonlinearity $\times$ $\times$ $\times$ Orthogonality $\times$ $\times$ $\times$ Large-scale monitoring $\times$ $\times$ $\times$ $\times$ $\times$ Adaptive Weight $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ Distribution monitoring $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
The rest of this paper is organized as follows. Section 2 provides a brief review of LSTM and the quantile-based multivariate CUSUM method, followed by a detailed introduction to the proposed MOLA and the adaptive W-BF frameworks. Section 3 discusses the detailed steps involved in offline training and online monitoring of the proposed process monitoring framework. Next, in Section 4, we showcase the performance of our proposed methodology in the benchmark problem of the Tennessee Eastman Process (TEP), demonstrating its outstanding fault detection speed and accuracy compared to benchmark methods. To conclude, we summarize all results and learnings and discuss potential improvements for future research in Section 5.
@@ 2.1 LSTM
The recurrent neural network (RNN) is a class of neural network architectures specifically designed for time series modeling and prediction. This makes RNN-based methods particularly attractive in capturing the dynamic features of industrial process data [28]. However, traditional RNNs are prone to the issues of gradient explosion and gradient vanishing when working with long-term time series [29]. To overcome these issues, LSTM, an advanced RNN that uses “gates” to capture both long-term and short-term memory, is introduced, featuring a unique structural unit at its core [30]. As illustrated in Figure 2, an LSTM unit contains three crucial control gates: the forget gate $f(t)$, the input gate $i(t)$, and the output gate $o(t)$. These gates work collectively to ensure that essential information is consistently retained while the less important information is discarded. Therefore, LSTM achieves superior performance in capturing complex temporal dynamics embedded in chemical process systems. For more detailed information and mathematical descriptions about LSTM, readers are encouraged to refer to [31].
@@ 2.2 OLAE
The LSTM AE is a specialized type of AE that seamlessly integrates LSTM with AE. This hybrid architecture enables efficient encoding and decoding of temporal sequences while capturing long-term dynamic features and dependencies within individual data streams at the same time. Similar to AE, the LSTM AE typically uses the reconstruction error as the loss function for model training. The Mean Squared Error (MSE) is one of the most widely adopted reconstruction error formulations:
@@
Figure 3: Our proposed OLAE architecture consists of an LSTM encoder, an LSTM decoder, and a fully connected (FC) layer that leverages orthogonality. The FC layer is denoted as $C=\sigma(Wh+b)$, where $h$ and $C$ represent the output of the encoder and fully connected (FC) layer, respectively. Here, $b$ and $\sigma$ are the bias term and nonlinear activation function of the FC layer, respectively.
Therefore, in the OLAE architecture as shown in Figure 3, we define an orthogonality-based loss function in Equation (2) to constrain the latent space output to generate non-redundant orthogonal latent features:
@@
where $W=[w_{1},w_{2},\dots,w_{m}]^{T}$ is the weight matrix of the FC layer. The loss function $\operatorname{Loss}_{\perp}$ consists of two components. The primary purpose of having the first component is to drive the inner products between the weight vectors of neurons in the FC layer towards zero, which indicates that the linear projection features from the encoder layer to the FC layer are mutually orthogonal. This component can also be viewed as a regularization term, which drives the weights of the FC layer to smaller values during model training and thus effectively mitigates overfitting issues. Meanwhile, the second term ensures that the extracted latent features remain as mutually independent as possible even after undergoing nonlinear transformation ($\sigma$).
@@ 2.3 Monitoring statistics
Overall, we monitor the two-sided statistic $W_{x}(t)=\max(W_{x}^{-}(t),W_{x}^{+}(t))$ for both upward and downward mean shifts. An alarm is raised (i.e., an anomaly is detected) when the monitoring statistic, $\sum_{(x)=1}^{r}W_{(x)}(t)$, defined as the sum of the largest $r$ local statistics $W_{x}$ at each time $t$, exceeds a threshold $h$ that is related to the pre-specified false alarm rate (e.g., 0.0027 for the typical $3\sigma$-limit) [34]. The corresponding stopping time $T$ is thus:
$T=\inf\left\{t>0:\sum_{(x)=1}^{r}W_{(x)}(t)\geq h\right\}.$ (8)
More information about the theory and application of quantile-based multivariate CUSUM, including detailed derivations of the mathematical formulations above, can be found in Ye and Liu [26].
@@ 2.4 Adaptive weight-based Bayesian fusion strategy
Here, we describe our multi-block monitoring framework for large-scale, complex industrial systems. Each block is monitored by two metrics (resp. $T^{2}$ and $W$) using two approaches (resp. Hotelling’s $T^{2}$ and quantile-based CUSUM). We adopt a Bayesian data fusion framework to aggregate the two monitoring results into a single monitoring metric in each block, as well as to aggregate all block-level metrics into a single plant-wide fault index (PFI). Bayesian fusion has shown remarkable robustness and capabilities in integrating information from diverse sources. It leverages prior knowledge and real-time measurements to compute posterior probabilities using Bayes’ theorem. In this work, we extend the classic Bayesian fusion methodology and propose an adaptive weight-based Bayesian fusion (W-BF) method. The idea is to dynamically adjust fusion weights based on the relative ranking of the current monitoring statistics from each block. Under this framework, the probability of sample $X_{i}^{n}(t)$ of block $n$ and monitoring metric $i$ in normal and fault conditions are given by:
@@
$B^{n}(t)=\frac{\sum_{i=1}^{2}w_{i}^{n}(t)\mathbb{P}_{i}^{n}(X_{i}^{n}(t)|F)\mathbb{P}^{n}_{i}(F|X_{i}^{n}(t))}{\sum_{i=1}^{2}w_{i}^{n}(t)\mathbb{P}_{i}^{n}(X_{i}^{n}(t)|F)},$ (11)
@@
$\operatorname{PFI}(t)=\frac{\sum_{n=1}^{N}w_{n}(t)\mathbb{P}^{n}(B^{n}(t)|F)\mathbb{P}^{n}(F|B^{n}(t))}{{\textstyle\sum_{n=1}^{N}}w_{n}(t)\mathbb{P}^{n}(B^{n}(t)|F)};$ (14)